OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (7, 4); Slope: 6 y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

B

OpenStudy (anonymous):

y-4 = 6( x-7) because the formula is y-y1 = m( x-x1)

OpenStudy (anonymous):

Thanks :)